Solve for y
x= W+Y+X /5

Answer by Nate(3500)

x= W+Y+X /5 (original equation)
x-w=y+x/5 (subtract w)
5x-5w=5y+x (multiply by 5 to simple the matters)
4x-5w=5y (subtract x)
(4/5)x-w=y (divide by 5 to get y by itself)